A NEW POTENTIAL ENERGY SURFACE FOR H_{2}–N_{2}O AND PIMC SIMULATION PROBING SUPER-FLUIDITY AND VIBRATIONAL FREQUENCY SHIFTS IN DOPED para-H_{2} CLUSTERS

\underline{LECHENG WANG}, ROBERT J LE ROY AND PIERRE-NICHOLAS ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada; DAIQIAN XIE, School of Chemistry and Chemical Engineering, Nanjing University, Nanjing, Jiangsu, China.

The existence of superfluidity of para-hydrogen clusters doped by the linear molecules^{a}^{,}^{b}^{,}^{c} such as CO, CO_{2} and OCS stimulated our theoretical study of (para-H_{2})_{N}–N_{2}O clusters. A new 6D ab initio PES for the H_{2}–N_{2}O dimer which explicitly included the symmetric and asymmetric vibrational coordinates Q_{1} and Q_{3} of N_{2}O was constructed. Fourdimensional global intermolecular PESs were then obtained by fitting the vibrational averaged interactions energies for ν_{3}(N_{2}O)=0 and 1 to the Morse/Long Range(MLR) analytical form with theoretically fixed long-range parameters. Using the adiabatic hindered-rotor approximation, effective two-dimensional PESs for the para-H_{2}–N_{2}O dimer were then generated. Predictions of the infrared spectra of para-H_{2}–N_{2}O based on these PESs have been in good agreement with the experimental observations.^{d} Based on these surfaces, predictions of structural properties, vibrational band origin shifts, rotational dynamics and superfluidity of (para-H_{2})_{N}–N_{2}O clusters have been generated using bosonic PIMC simulation methods. The evolution of the calculated shifts agreed reasonably with that for the experimental observations.^{e} Reduction of the effective moment of inertia is predicted to occur when the dopant is partially surrounded by the para-H_{2} solvent, which marked the onset of molecular superfluidity in para-H_{2}. Results obtained using our surfaces and PIMC algorithm will be presented.

a H. Li, R. J. Le Roy, P.-N. Roy and A.R.W. McKellar, Phys. Rev. Lett. 105, 133401 (2010).

b S. Grebenev, B. G. Sartakov, J. P. Toennies and A. F. Vilesov, J. Chem. Phys. 132, 064501 (2010).

cP. L. Raston, W. Jager, H. Li, R. J. Le Roy, and P.-N. Roy , Phys. Rev. Lett. 108, 253402 (2012).

d J. Tang and A.R.W. McKellar, J. Chem. Phys. 117, 8308 (2002).

e J. Tang and A.R.W. McKellar, J. Chem. Phys. 123, 114314 (2005).